The mean of the first five terms of an arithmetic sequence is 27, and the mean of the first eight terms of the same sequence is 39. What is the absolute difference between the first and second terms of this sequence?

Answer by mananth(16946)

Mean of First term of an AP is 27
Sum = 5*27 = 135
S5= 135
similarly S8 = 39*8= 312
135 = 5/2( 2a+(5-1)d)
270=5( 2a+4d)
54 =2a+4d
/2
27= a+2d..........(1)
sum of eighth term
312 = 8/2(2a+(8-1)d)
78 = 2a+7d..........(2)
solve (1) & (2)
we get d=8
and a = 11
I term = 11
difference =8